OpenStudy (anonymous):

Some one help me PLEASE Find the quotient 2 1/2 divided by 1/4

OpenStudy (solomonzelman):

Just like\[\huge\color{blue}{ a \frac{b}{c} =\frac{a \times c + b}{c} }\]so to \[\huge\color{blue}{ 2 \frac{1}{2} =\frac{2 \times 2 + 1}{2} }\]

OpenStudy (solomonzelman):

So we know that\[\huge\color{blue}{ 2 \frac{1}{2} = }\huge\color{blue}{ \frac{5}{2} }\]

OpenStudy (solomonzelman):

No lets divide, the problem is now saying, \[\huge\color{red}{ \frac{5}{2} \div }\huge\color{red}{ \frac{1}{4} }\]

OpenStudy (anonymous):

SO is it 5/8?

OpenStudy (anonymous):

well is the answer 5/8?

OpenStudy (solomonzelman):

when you divide you flip the sign and multiply. \[\huge\color{red}{ \frac{5}{2} }\times \huge\color{red}{ \frac{4}{1} }\]

OpenStudy (solomonzelman):

I should have said. when you divide by a fraction, you flip the fraction you are dividing be, and multiply instead. see what I did?

OpenStudy (anonymous):

oh ok so is the answer 20/2 but that does not sound right

OpenStudy (solomonzelman):

Yes, because you can reduce\[\huge\color{red}{ \frac{20}{2} }\]

OpenStudy (anonymous):

so is the answer 10? thats what i got

OpenStudy (solomonzelman):

Yes!

OpenStudy (anonymous):

Thanks sooooooooooo much!!!!!!!!!!!!!!

OpenStudy (solomonzelman):

You welcome!

OpenStudy (anonymous):

:)